# A stochastic approach in physics exercises of mathematics education

### v1

@@ 3 A stochastic approach in statics
In this section, we demonstrate our method for incorporating a stochastic point of view into physics exercises of mathematics education by presenting exercises for statics. In them, we randomize the geometry of the static problem, for example, the shape of a plate or the position of some of the pinning points of a plate. To avoid technicalities, we choose uniform distributions as the distributions of the random inputs. Altogether, we present three exercises in this section.
@@
A square piece of homogeneous sheet metal of side length 1 m and mass $m$ kg is bent in the following way: on one side, at a point chosen at a distance of $\frac{1}{4}$ from one of the vertices, a perpendicular is set on that side, and we bend the metal sheet along this perpendicular in a way that the angle between the two pieces is uniformly distributed on the interval $(0,\frac{\pi}{2})$. We place the bent plate with its smaller part on a horizontal surface, and then release it.
@@
Figure 6:
@@
$B:=$the perpendicular projection of S1 does not fall on the section $\Big{\{}\text{the perpendicular projection of $S_{1}$ does not fall on the section $TR$}\Big{\}}$
@@
$B$ =$=\left\{\cos(\alpha)>\frac{2}{3}\right\}\cap\left\{\frac{1}{24}\geqslant\frac{3}{8}\cos(\alpha)-\frac{1}{4}\right\}=\left\{\cos(\alpha)>\frac{2}{3}\right\}\cap\left\{\cos(\alpha)\leqslant\frac{7}{9}\right\}$
=$=\left\{\frac{2}{3}<\cos(\alpha)\leqslant\frac{7}{9}\right\}.$
[…]
@@
({the bent plate does not fall $\operatorname{\mathbb{P}}(\{\text{the bent plate does not fall over}\})=\operatorname{\mathbb{P}}\left(\alpha_{0}\leqslant\alpha<\frac{\pi}{2}\right)$
@@
Figure 7:
Solution. (i). Let $C$ and $\xi$ denote the point corresponding to the roller support and the length of the section $SC$, respectively. Due to the assumptions, $\xi$ is a random variable having uniform distribution on the interval $[0,\frac{1}{2}]$. The reaction forces at the points $A$ and $C$ are denoted by ${\boldsymbol{F}}_{A}=[F_{A,x},F_{A,y}]^{\top}$ and ${\boldsymbol{F}}_{C}=[F_{C,x},F_{C,y}]^{\top}$, respectively. Furthermore, let ${\boldsymbol{F}}_{S}:=[0,-mg]^{\top}$. Note also that the direction of rotation is positive (i.e., counterclockwise) on Figure 7.
@@
where we used Newton-Leibniz formula and that a primitive function of the function $[0,\frac{1}{2}]\ni x\mapsto 2/(2x+1)$ is $[0,\frac{1}{2}]\ni x\mapsto\ln(2x+1)$.
@@
The plate with the dimensions given in Figure 8 is fixed by means of a pin support at point $A$ and a rope at point $B$, and is subjected to concentrated forces ${\boldsymbol{F}}_{1}$ and ${\boldsymbol{F}}_{2}$, respectively. Assume that the mass of the plate can be neglected, and the magnitude of the forces ${\boldsymbol{F}}_{1}$ and ${\boldsymbol{F}}_{2}$ is 300300300\,300N and 100100100\,100N, respectively. Further, we suppose that the distribution of $b$ is uniform distribution on the interval $(0\,\mathrm{m},\sqrt{3}\,\mathrm{m})$.
Figure 8:
@@
[…]
@@
=$=\begin{cases}0&\text{if $x\leqslant 400\sqrt{3}+\frac{100}{3}\sqrt{2}$,}\\ \frac{x-400\sqrt{3}+\frac{100}{3}\sqrt{2}}{200\sqrt{3}}&\text{if $x\in\left(400\sqrt{3}-\frac{100}{3}\sqrt{2},400\sqrt{3}-\frac{100}{3}\sqrt{2}+600\right)$,}\\ 1&\text{if $x\geqslant 400\sqrt{3}-\frac{100}{3}\sqrt{2}+600$.}\end{cases}$
@@ 4 Concluding remarks
Our method for incorporating a stochastic point of view into physics exercises of mathematics education may be used in several different ways by high school students and teachers and by first-year university students. The core of the method is the randomization of some inputs, the system model used does not differ from what we would use in the deterministic approach. We demonstrate our method of randomization of some inputs by considering exercises from the theory of projectile motion and statics. Students may be asked to create similar exercises, for other fields of physics as well. Our stochastic perspective can help the students recognize the fact that many phenomena in our daily lives have a random nature, and therefore it can be useful to get in touch with stochastic thinking and tools in order to be able to understand these phenomena much better. Computer simulations using Monte Carlo methods might also be assigned to our elaborate examples in order to confirm the solutions empirically.
@@ Appendix A Basic knowledge of probability
The cumulative distribution function of a random variable $\xi:\Omega\to\mathbb{R}$ is defined by $F_{\xi}:\mathbb{R}\to\mathbb{R}$, $F_{\xi}(x):=\operatorname{\mathbb{P}}(\xi<x)$, $x\in\mathbb{R}$. A random variable is called discrete if its range $\xi(\Omega)$ is countable. By the (probability) distribution of a discrete random variable $\xi$ having range $\xi(\Omega)=:\{x_{i}:i\in\mathbb{N}\}$, we mean the sequence $(\operatorname{\mathbb{P}}(\xi=x_{i}))_{i\in\mathbb{N}}$. A random variable $\xi:\Omega\to\mathbb{R}$ is called absolutely continuous if there exists a (Borel measurable) non-negative function $f_{\xi}:\mathbb{R}\to\mathbb{R}$ such that $F_{\xi}(x)=\int_{-\infty}^{x}f_{\xi}(t)\,\mathrm{d}t$, $x\in\mathbb{R}$. The expected value (expectation) and of a discrete random variable $\xi$ having range $\xi(\Omega)=:\{x_{i}:i\in\mathbb{N}\}$ is defined by $\operatorname{\mathbb{E}}(\xi):=\sum_{i=1}^{\infty}x_{i}\operatorname{\mathbb{P}}(\xi=x_{i})$, provided that $\sum_{i=1}^{\infty}|x_{i}|\operatorname{\mathbb{P}}(\xi=x_{i})<\infty$. The expected value (expectation) of an absolutely continuous random variable $\xi:\Omega\to\mathbb{R}$ with a density function $f_{\xi}$ is defined by $\operatorname{\mathbb{E}}(\xi):=\int_{-\infty}^{\infty}xf_{\xi}(x)\,\mathrm{d}x$, provided that $\int_{-\infty}^{\infty}|x|f_{\xi}(x)\,\mathrm{d}x<\infty$. The variance and standard deviation of a random variable $\xi$ is defined by $\operatorname{\mathbb{D}}^{2}(\xi):=\operatorname{\mathbb{E}}((\xi-\operatorname{\mathbb{E}}(\xi))^{2})$ and $\operatorname{\mathbb{D}}(\xi):=\sqrt{\operatorname{\mathbb{E}}((\xi-\operatorname{\mathbb{E}}(\xi))^{2})}$, respectively, provided that $\operatorname{\mathbb{E}}(\xi^{2})<\infty$. It is know that $\operatorname{\mathbb{D}}^{2}(\xi)=\operatorname{\mathbb{E}}(\xi^{2})-(\operatorname{\mathbb{E}}(\xi))^{2}$. The relative standard deviation (also known as coefficient of variation) of a positive random variable $\xi$ having finite second moment (i.e., $\operatorname{\mathbb{E}}(\xi^{2})<\infty$) is given by $\operatorname{RSD}(\xi):=\operatorname{\mathbb{D}}(\xi)/\operatorname{\mathbb{E}}(\xi)$. Roughly speaking, the higher relative standard deviation of a nonnegative random variable $\xi$, the greater the level of dispersion of the values of $\xi$ around its mean $\operatorname{\mathbb{E}}(\xi)$. If $(\xi,\eta)$ is a $2$-dimensional random vector such that $\operatorname{\mathbb{E}}(\xi)$ and $\operatorname{\mathbb{E}}(\eta)$ are defined, then the expectation vector of $(\xi,\eta)$ is given by $\operatorname{\mathbb{E}}(\xi,\eta):=(\operatorname{\mathbb{E}}(\xi),\operatorname{\mathbb{E}}(\eta))$. If $\xi:\Omega\to\mathbb{R}$ is a random variable such that $\operatorname{\mathbb{E}}(\xi)$ is defined and $a,b\in\mathbb{R}$, then $\operatorname{\mathbb{E}}(a\xi+b)$ is defined as well and
@@ Acknowledgements
We would like to thank Gábor Drótos, Janka Vértesi and Tamás Vértesi for their helpful comments and suggestions. We are grateful to Renáta Vas for helping us with the language.

### v2

@@ 3 A stochastic approach in statics
In this section, we demonstrate our method for incorporating a stochastic point of view into physics exercises of mathematics education by presenting exercises for statics. In them, we randomize the geometry of the static problem, for example, the shape of a plate or the position of some of the pinning points of a plate. To avoid technicalities, we choose uniform distributions as the distributions of the random inputs. Altogether, we present three exercises in this section. These exercises highlight the importance of probabilistic tools in static problems, they can make abstract concepts of probability theory more tangible.
@@
A square piece of homogeneous sheet metal of side length 1 m and mass $m$ kg is bent in the following way: on one side, at a point chosen at a distance of $\frac{1}{4}$ from one of the vertices, a perpendicular is set on that side, and we bend the metal sheet along this perpendicular in a way that the angle between the two pieces is uniformly distributed on the interval $(0,\frac{\pi}{2})$. We place the bent plate with its smaller part on a horizontal surface (see Figure 6), and then release it.
@@
Figure 6: a planar section of the bent plate for Exercise 3.1 and its solution.
@@
(i). Then $\{\text{the bent plate does not fall over}\}=A\cup B$, where the events $A$ and $B$ are defined by
@@
[…]
@@
$B$ $=\left\{\cos(\alpha)>\frac{2}{3}\right\}\cap\left\{\frac{1}{24}\geqslant\frac{3}{8}\cos(\alpha)-\frac{1}{4}\right\}=\left\{\cos(\alpha)>\frac{2}{3}\right\}\cap\left\{\cos(\alpha)\leqslant\frac{7}{9}\right\}$
$=\left\{\frac{2}{3}<\cos(\alpha)\leqslant\frac{7}{9}\right\}.$
Hence $\{\text{the bent plate does not fall over}\}=\{\cos(\alpha)\leqslant\frac{7}{9}\}$. Here the inequality $\cos(\alpha)\leqslant\frac{7}{9}$, $\alpha\in(0,\frac{\pi}{2})$ holds if and only if $\alpha\in[\alpha_{0},\frac{\pi}{2})$, where $\alpha_{0}\in(0,\frac{\pi}{2})$ is the unique value for which $\cos(\alpha_{0})=\frac{7}{9}$, i.e., $\alpha_{0}=\arccos\left(\frac{7}{9}\right)\approx 0.6797$. Consequently, we have
$\{\text{the bent plate does not fall over}\}=A\cup B=\left\{\alpha_{0}\leqslant\alpha<\frac{\pi}{2}\right\},$
@@
$\operatorname{\mathbb{P}}(\{\text{the bent plate does not fall over}\})=\operatorname{\mathbb{P}}\left(\alpha_{0}\leqslant\alpha<\frac{\pi}{2}\right)$
@@
Figure 7: a prismatic beam fixed by means of a pin support at point $A$ and a roller support at point $C$ for Exercise 3.2 and its solution.
Solution. (i). Let $C$ and $\xi$ denote the point corresponding to the roller support and the length of the section $SC$, respectively. Due to the assumptions, $\xi$ is a random variable having uniform distribution on the interval $0,\frac{1}{2}]$. The reaction forces at the points $A$ and $C$ are denoted by ${\boldsymbol{F}}_{A}=[F_{A,x},F_{A,y}]^{\top}$ and ${\boldsymbol{F}}_{C}=[F_{C,x},F_{C,y}]^{\top}$, respectively. Furthermore, let ${\boldsymbol{F}}_{S}:=[0,-mg]^{\top}$. Note also that the direction of rotation is positive (i.e., counterclockwise) on Figure 7.
@@
where we used Newton-Leibniz formula and that a primitive function of the function $0,\frac{1}{2}]\ni x\mapsto 2/(2x+1)$ is $0,\frac{1}{2}]\ni x\mapsto\ln(2x+1)$.
@@
The plate with the dimensions given in Figure 8 is fixed by means of a pin support at point $A$ and a rope at point $B$, and is subjected to concentrated forces ${\boldsymbol{F}}_{1}$ and ${\boldsymbol{F}}_{2}$, respectively. Assume that the mass of the plate can be neglected, and the magnitude of the forces ${\boldsymbol{F}}_{1}$ and ${\boldsymbol{F}}_{2}$ is 300300\,N and 100100\,N, respectively. Further, we suppose that the distribution of $b$ is uniform distribution on the interval $(0\,\mathrm{m},\sqrt{3}\,\mathrm{m})$.
Figure 8: a plate fixed by means of a pin support at point $A$ and a rope at point $B$, and is subjected to concentrated forces ${\boldsymbol{F}}_{1}$ and ${\boldsymbol{F}}_{2}$, respectively for Exercise 3.3 and its solution.
@@
We take the direction of rotation as positive (i.e., counterclockwise) on Figure 8. Hence the moments $\underset{A}{M_{1}}$, $\underset{A}{M_{2}}$ and $\underset{A}{M_{B}}$ of the forces ${\boldsymbol{F}}_{1}$, ${\boldsymbol{F}}_{2}$ and ${\boldsymbol{F}}_{B}$ with respect to the point $A$ are
$\underset{A}{M_{1}}:=\|{\boldsymbol{F}}_{1}\|\cdot\cos\left(\frac{30}{360}\cdot 2\pi\right)\cdot(2+b)=300\frac{\sqrt{3}}{2}(2+b)=150\sqrt{3}(2+b),$
$\underset{A}{M_{2}}:=\|{\boldsymbol{F}}_{2}\|\cdot\cos\left(\frac{45}{360}\right)\cdot 1.5-\|{\boldsymbol{F}}_{2}\|\cdot\sin\left(\frac{45}{360}\right)\cdot 2=150\frac{\sqrt{2}}{2}-100\sqrt{2}=-25\sqrt{2},$
and
$\underset{A}{M_{B}}:=-\|{\boldsymbol{F}}_{B}\|\cdot\cos\left(\frac{60}{360}\cdot 2\pi\right)\cdot 1.5=-\|{\boldsymbol{F}}_{B}\|\cdot\frac{1}{2}\cdot 1.5=-\frac{3}{4}\|{\boldsymbol{F}}_{B}\|.$
@@
$=\begin{cases}0&\text{if $x\leqslant 400\sqrt{3}+\frac{100}{3}\sqrt{2}$,}\\ \frac{x-400\sqrt{3}+\frac{100}{3}\sqrt{2}}{200\sqrt{3}}&\text{if $x\in\left(400\sqrt{3}-\frac{100}{3}\sqrt{2},400\sqrt{3}-\frac{100}{3}\sqrt{2}+600\right)$,}\\ 1&\text{if $x\geqslant 400\sqrt{3}-\frac{100}{3}\sqrt{2}+600$.}\end{cases}$
@@
Regarding didactic conclusions, we can say that our exercises can draw the learners’ attention to the importance of the random positioning of the structural elements and its consequences in static problems.
@@ 4 Concluding remarks
Our method for incorporating a stochastic point of view into physics exercises of mathematics education may be used by students and teachers familiar with elementary probability theory and mechanics for understanding some basic concepts of stochastic mechanics. The core of the method is the randomization of some inputs, the system model used does not differ from what we would use in the deterministic approach. We demonstrate our method of randomization of some inputs by considering exercises from the theory of projectile motion and statics. STEM students may be asked to create similar exercises, for other fields of physics as well. Our stochastic perspective can help the students recognize the fact that many phenomena in our daily lives have a random nature, and therefore it can be useful to get in touch with stochastic thinking and tools in order to be able to understand these phenomena much better. Computer simulations using Monte Carlo methods might also be assigned to our elaborate examples in order to confirm the solutions empirically.
@@ Appendix A Basic knowledge of probability
The cumulative distribution function of a random variable $\xi:\Omega\to\mathbb{R}$ is defined by $F_{\xi}:\mathbb{R}\to\mathbb{R}$, $F_{\xi}(x):=\operatorname{\mathbb{P}}(\xi<x)$, $x\in\mathbb{R}$. A random variable is called discrete if its range $\xi(\Omega)$ is countable. By the (probability) distribution of a discrete random variable $\xi$ having range $\xi(\Omega)=:\{x_{i}:i\in\mathbb{N}\}$, we mean the sequence $(\operatorname{\mathbb{P}}(\xi=x_{i}))_{i\in\mathbb{N}}$. A random variable $\xi:\Omega\to\mathbb{R}$ is called absolutely continuous if there exists a (Borel measurable) non-negative function $f_{\xi}:\mathbb{R}\to\mathbb{R}$ such that $F_{\xi}(x)=\int_{-\infty}^{x}f_{\xi}(t)\,\mathrm{d}t$, $x\in\mathbb{R}$. The expected value (expectation) of a discrete random variable $\xi$ having range $\xi(\Omega)=:\{x_{i}:i\in\mathbb{N}\}$ is defined by $\operatorname{\mathbb{E}}(\xi):=\sum_{i=1}^{\infty}x_{i}\operatorname{\mathbb{P}}(\xi=x_{i})$, provided that $\sum_{i=1}^{\infty}|x_{i}|\operatorname{\mathbb{P}}(\xi=x_{i})<\infty$. The expected value of a discrete random variable having finite range is nothing else but the weighted average of the possible outcomes of the random variable, where the weights are the probabilities that the outcomes will occur. The expected value (expectation) of an absolutely continuous random variable $\xi:\Omega\to\mathbb{R}$ with a density function $f_{\xi}$ is defined by $\operatorname{\mathbb{E}}(\xi):=\int_{-\infty}^{\infty}xf_{\xi}(x)\,\mathrm{d}x$, provided that $\int_{-\infty}^{\infty}|x|f_{\xi}(x)\,\mathrm{d}x<\infty$. The variance and standard deviation of a random variable $\xi$ is defined by $\operatorname{\mathbb{D}}^{2}(\xi):=\operatorname{\mathbb{E}}((\xi-\operatorname{\mathbb{E}}(\xi))^{2})$ and $\operatorname{\mathbb{D}}(\xi):=\sqrt{\operatorname{\mathbb{E}}((\xi-\operatorname{\mathbb{E}}(\xi))^{2})}$, respectively, provided that $\operatorname{\mathbb{E}}(\xi^{2})<\infty$. It is know that $\operatorname{\mathbb{D}}^{2}(\xi)=\operatorname{\mathbb{E}}(\xi^{2})-(\operatorname{\mathbb{E}}(\xi))^{2}$. The standard deviation of a random variable can be interpreted as a measure of the amount of spread of the values of the random variable around its expected value. A higher standard deviation indicates a greater spread, meaning the random variable is more likely to take on values far from its expected value. The relative standard deviation (also known as coefficient of variation) of a positive random variable $\xi$ having finite second moment (i.e., $\operatorname{\mathbb{E}}(\xi^{2})<\infty$) is given by $\operatorname{RSD}(\xi):=\operatorname{\mathbb{D}}(\xi)/\operatorname{\mathbb{E}}(\xi)$. Roughly speaking, the higher relative standard deviation of a nonnegative random variable $\xi$, the greater the level of dispersion of the values of $\xi$ around its mean $\operatorname{\mathbb{E}}(\xi)$. If $(\xi,\eta)$ is a $2$-dimensional random vector such that $\operatorname{\mathbb{E}}(\xi)$ and $\operatorname{\mathbb{E}}(\eta)$ are defined, then the expectation vector of $(\xi,\eta)$ is given by $\operatorname{\mathbb{E}}(\xi,\eta):=(\operatorname{\mathbb{E}}(\xi),\operatorname{\mathbb{E}}(\eta))$. If $\xi:\Omega\to\mathbb{R}$ is a random variable such that $\operatorname{\mathbb{E}}(\xi)$ is defined and $a,b\in\mathbb{R}$, then $\operatorname{\mathbb{E}}(a\xi+b)$ is defined as well and
@@ Appendix B Basic knowledge of mechanics
Appendix C Alternative solution to part (ii) of Exercise 2.1
In this appendix, we determine the probabilities of the events $R$ and $S$ introduced in the solution of part (ii) of Exercise 2.1 using the law of total expectation. This is an alternative solution, since in the solution of part (ii) of Exercise 2.1, we used the method of geometric probabilities.
First, we determine the probability $\operatorname{\mathbb{P}}(R)$. Since $v$ is uniformly distributed on the interval $[5,15]$, its density function takes the form
where
since $50g/r^{2}\leqslant 50g/25=19.62$. Consequently,
Next, we determine the probability $\operatorname{\mathbb{P}}(S)$. Using the law of total expectation (see (A.4)) and the independence of $h$ and $v$, we have that
Note that we calculated the same values for $\operatorname{\mathbb{P}}(R)$ and $\operatorname{\mathbb{P}}(S)$ as in the solution of part (ii) of Exercise 2.1, that was expected.
@@ Acknowledgements
We would like to thank Gábor Drótos, Janka Vértesi and Tamás Vértesi for their helpful comments and suggestions. We are grateful to Renáta Vas for helping us with the language. We would like to thank the three referees for their comments that helped us to improve the paper.